Order: Vitamin K (Aquamephyton) 1mg IM x1 dose

Weight: 4kg

Supply: 10mg/ml

Literature: prophylaxis and treatment: 0.5 to 1 mg/dose IM x1

How many ml will you give? _____Is this a safe dose? YES or NO

2. Order: amoxicillin/clavulante (augmentin) 125 mg po Q8h

Weight: 1 year old, 10 kg

Supply: 125mg/ 5ml

Literature: safe dose amoxicillin/clavulante: 20 to 40 mg/kg/24 hours given q8h po

How many ml will you give? _____Is this a safe dose? YES or NO

3. Order: benzathine penicillin 500,000 units IM x1 dose

Weight: 10 month old, 10 kg

Supply: vial labeled 600,000 units/ml

Literature: safe dose 50,000 units/kg x 1 IM. Maximum, 2.4 milliunits

How many ml will you give? _____Is this a safe dose? YES or NO
Answer #1

ANSWER.

1) Calculation of amount(ml) of drug will you give.

      Formula

        Volume(ml) of drug required = Desired dose/Dose in hand * Quantity.

    Here

      Desired dose (D)= 1 mg

     Dose in hand (H) = 10 mg

       Quantity (Q) = 1 ml

As per above formula

       Volume(ml) of drug required = D/H * Q

                                                     = 1/10 * 1

                                                     = 0.1 ml

Dose for prophylaxis and treatment is 0.5 to 1 mg,

      So this dose is safe for this child. (YES)

2) Calculation of volume(ml) of drug required.

      Formula

        Volume(ml) = Desired dose/Dose in hand * Quantity.

     Here

        Desired dose (D) = 125 mg.

        Dose in hand (H) = 125 mg

            Quantity (Q) = 5 ml

As per above formula

        Volume(ml) of drug required = D/H * Q

                                                      = 125/125 * 5

                                                      = 5 ml.

   Literature safe dose = 20 to 40 mg/kg/24 hrs.

   Dose suitable for a 10 kg child = 20 * 10 to 40 * 10

                                                    = 200 mg to 400 mg per day.

      Ordered dose = 125 mg ( q8hr or three times a day)

                             = 125 * 3

                             = 375 mg.

   * So the ordered dose is safe for this child (YES)

3) Calculation of volume(ml) of drug required.

        Formula

          Volume of drug required = Desired dose /Dose in hand * Quantity

      Here

         Desired dose (D) = 500000 units.

         Dose in hand (H) = 600000 units.

            Quantity (Q) = 1 ml

   As per above formula

       Volume(ml) of drug required = D/H * Q

                                                     = 500000 / 600000 * 1

                                                     = 0.8 *1

                                                     = 0.8 ml

        Safe dose = 50000 units/kg.(Maximum 2.4 milliunits).

        Dose required for a 10 kg child = 50000 * 10

                                                           = 500000 units.

       Ordered dose = 500000 units.

   * So the ordered dose is safe for this child (YES).